OpenStudy (anonymous):

find exact value of tan(255) using half angle identity for tangent. Half angle identity is tan(a/2) but i want to approach this problem with finding two angles that I know like 225 degrees + 30 degrees. Any ideas?

OpenStudy (solomonzelman):

Do the half angle man !! \(\large\color{blue}{ \bf tan(255)=tan(510/2)=±\sqrt{ \frac{1-cos(510)}{1+cos(510)} } }\) \(\large\color{blue}{ \bf ±\sqrt{ \frac{1-cos(510)}{1+cos(510)} } =\frac{1-cos(510)}{sin(510)} }\)

OpenStudy (solomonzelman):

cos(510)=cos(480+30)=cos(480)cos(30)-sin(480)sin(30) = ? sin(510)=sin(480+30)=sin(480)cos(30)+cos(480)sin(30) = ? go for it....

OpenStudy (anonymous):

Awesome I didn't see that 510/2 = 255 which puts me at cords. \[(-\sqrt{3})/2, 1/2)\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

\[(-\sqrt{3}/2, 1/2)\] somehow an extra ")" got in there.

OpenStudy (solomonzelman):

Looks right, I just don;t know the values of sin, cos 480 I know that sin(30)=1/2 and cos(30=√3 / 2 (each time I forget the 30º sine and cosine, but then draw a 30-60-90 triangle and apply sohcahtoa) .... have a good night....

OpenStudy (anonymous):

same to you.